OpenStudy (anonymous):

solve system y=-3x^2+x+4 y=-2x-2

OpenStudy (anonymous):

If they both equal y that means they equal each other. So you can get x with: \[-3x^2 + x + 4 = -2x - 2\]

OpenStudy (anonymous):

work?

OpenStudy (anonymous):

-3x^2+x+4=-2x-2 -3x^2+4=-3x-2 -3x^2=-3x-6 -3x^2+3x+6=0 (-3x+6)(x-3)=0 -3x+6= 0 x-3=0 -3x=-6 x=3 x=-6/-3=6/3 x=2